Suppose a meter stick made of steel and one made of invar are the same length at 0°C. What is their difference in length at 59.5°C? The coefficient of thermal expansion is 12 ✕ 10−6/°C for steel and 0.9 ✕ 10−6/°C for invar. mm

(b) Repeat the calculation for two 34.5-m-long surveyor's tapes. mm
Answer #1

a) Change in length is given by

\Delta L=L*\Delta T*(\alpha _{s}-\alpha _{i})    (1)

where L is the length = 1m , VI is the change in temperature=59.50C, \alpha _{S} is thermal expansion coefficient of steel =12x10-6/0C, \alpha _{i} is thermal expansion coefficient of ivar= 0.9 ✕ 10-6/°C

substitute

\Delta L=1*59.6*(12*10^{-6}-0.9*10^{-6})=6.6*10^{-4}m

b) here only change is length of the tape 34.5 m substitute in (1)

\Delta L=34.5*59.6*(12*10^{-6}-0.9*10^{-6})=6.6*10^{-4}m=0.02m